# Auditing Google’s AI Overviews and Featured Snippets: A Case Study on Baby Care and Pregnancy

### v1

@@ Introduction
Studies specifically auditing and evaluating generative AI search engines are limited, with few early explorations conducting evaluation via either fully manual annotation (Li and Aral 2025; Liu, Zhang, and Liang 2023; Hu et al. 2024) or hybrid methods combining manual effort with LLM-as-a-judge or machine learning classifiers (Venkit et al. 2024; Li and Sinnamon 2024). Most focus on specific answer quality properties (e.g., verifiability, sentiment, source category) or scenarios (e.g., adversarial attacks). For example, prior work (Liu, Zhang, and Liang 2023) examined verifiability through broad metrics (fluency, perceived utility) for measuring answer’s text generally and specific metrics (citation recall, citation precision) for checking answer’s reference, finding that generative AI responses often exhibit high fluency and perceived utility but frequently contain unsupported statements or inaccurate citations. Another study (Hu et al. 2024) evaluated generative search engines under adversarial input scenarios, demonstrating effectiveness in inducing incorrect responses via manipulative questions. Research in Li and Sinnamon (2024) examined sentiment in queries and answers, as well as source types, finding evidence of response sentiment bias driven by query sentiment, plus commercial and geographic biases in source categories. Research in Li and Aral (2025) specifically studied people’s trust in GenAI search engines compared to traditional ones via a preregistered, randomized experiment on a large U.S. representative sample.
Among these works, Venkit et al. (2024) conducted a more comprehensive evaluation spanning answer text, citations, sources, and user interface via hybrid methods including LLM-as-a-judge and manual evaluation (user interviews). They found that GenAI answers tend to use overtly confident language and align with biases implied in questions (“confirmation bias”). However, while confirmation bias wasn’t specifically studied during implementation in (Venkit et al. 2024), and (Li and Sinnamon 2024) did not control for specific topics or question types – factors known to affect question-answering systems (Li and Aral 2025; Zhao et al. 2019; Buck et al. 2017; Chu et al. 2020) – we investigate confirmation bias more rigorously by controlling query topic and question type (RQ4). We also integrate metrics from prior studies relevant to our research setting, which we elaborate on in the following methodology section. Given previous work’s caution about biases in LLM-as-a-judge evaluation (Ye et al. 2024) and the only moderate correlation between human annotation and LLM-as-a-judge (e.g., Pearson correlation of 0.62 in similar search settings (Venkit et al. 2024)), we opt for full manual evaluation for all metrics except source category in this study.
@@
For the first step, we selected the queries related to baby care and pregnancy from the 10.4 million distinct Bing search queries in the Open Resource for Click Analysis in Search (ORCAS) (which is a public click-based dataset associated with information retrieval tasks [^1] This search query dataset contains no personally identifiable information and has been filtered using a k-anonymity threshold to exclude infrequent queries. Offensive and pornographic content has been removed, meaning some extreme queries may be absent.(Craswell et al. 2020). Using this dataset allows us to include real search queries actually utilized by the users when seeking information or advice on baby care and pregnancy-related questions. For the first filtering step, we used the keywords “baby” and “babies” for the topic baby care and “pregnant” and “pregnancy” for the topic pregnancy. For the second step, we filtered for queries with a minimum length of 3 words, as 2-word queries (e.g., “baby teeth,” “baby eyes”) tend to be too vague. Additionally, we removed almost identical queries. To do this, we applied tokenization, stemming, and stopwords-removal to the queries, and then removed the queries that were duplicate after this pre-processing step. As a result, we obtained a total of 15,500 queries.
In the third step, we conducted further filtering of the queries to ensure we include only those related to human baby care and/or pregnancy. This was necessary to filter out queries related to (1) baby care/pregnancy of an animal or plant [^2] We filtered out any plant-related queries that might use pregnancy-related terminology.; (2) information about pregnancy of a celebrity, fictional character, or a specific person; (3) a name, song, or item that simply contains the words “baby” or “pregnancy” but is not related to seeking advice or information on pregnancy. For this step, we relied on annotation using gpt-4o-mini. Before using it on our full dataset, we validated the performance of the model on a subset of 50 manually annotated queries, and gpt-4o-mini classifications overlapped with our manual annotations 100% of the time. The exact prompt we used can be found in the Appendix. In total, after this step we retained 9516 queries relevant for questions about human baby care and/or pregnancy.
@@
  • Numeric Mismatches: Different, but specific ranges, times, measurements or likelihoods that don’t overlap and can create confusion; any numerical or temporal or likelihood differences or mismatch, following a definition in prior work (Li, Raheja, and Kumar 2023);
@@ Results
Query sentiment is also significantly (Chi-Square test $p<0.001$) associated with the frequency of appearance of FS. FS appeares nearly twice as often for negative (48.41%) compared to neutral (24.84%) queries. The difference is also statistically significant ($p=0.0265$) for AIO, with AIO appearing most often for negative (86.62%) and neutral (85.99%) queries, and slightly less often for positive ones (76.43%) (as shown in Figure 2) (b).

### v2

@@ Introduction
Studies specifically auditing and evaluating generative AI search engines are limited, with few early explorations conducting evaluation via either fully manual annotation (Li and Aral 2025; Liu, Zhang, and Liang 2023; Hu et al. 2024) or hybrid methods combining manual effort with LLM-as-a-judge or machine learning classifiers (Venkit et al. 2024; Li and Sinnamon 2024). Most focus on specific answer quality properties (e.g., verifiability, sentiment, source category) or scenarios (e.g., adversarial attacks). For example, prior work (Liu, Zhang, and Liang 2023) examined verifiability through broad metrics (fluency, perceived utility) for measuring answers’ text generally and specific metrics (citation recall, citation precision) for checking answers’ reference, finding that generative AI responses often exhibit high fluency and perceived utility but frequently contain unsupported statements or inaccurate citations. Another study (Hu et al. 2024) evaluated generative search engines under adversarial input scenarios, demonstrating effectiveness in inducing incorrect responses via manipulative questions. Research in Li and Sinnamon (2024) examined sentiment in queries and answers, as well as source types, finding evidence of response sentiment bias driven by query sentiment, plus commercial and geographic biases in source categories. Research in Li and Aral (2025) specifically studied people’s trust in GenAI search engines compared to traditional ones via a preregistered, randomized experiment on a large U.S. representative sample.
Among these works, Venkit et al. (2024) conducted a more comprehensive evaluation spanning answer text, citations, sources, and user interface via hybrid methods including LLM-as-a-judge and manual evaluation (user interviews). They found that GenAI answers tend to use overtly confident language and align with biases implied in questions (“confirmation bias”). However, while confirmation bias wasn’t specifically studied during implementation in (Venkit et al. 2024), and (Li and Sinnamon 2024) did not control for specific topics or question types – factors known to affect question-answering systems (Li and Aral 2025; Zhao et al. 2019; Buck et al. 2017; Chu et al. 2020) – we investigate confirmation bias more rigorously by controlling query topic and question type (RQ4). We also integrate metrics from prior studies relevant to our research setting, which we elaborate on in the following methodology section. Previous work has raised concerns about biases in LLM-as-a-judge evaluation (Ye et al. 2024) and documented only moderate correlation between human annotation and LLM-as-a-judge (e.g., Pearson correlation of 0.62 in similar search settings (Venkit et al. 2024)), while recent evidence has documented systematic risks and unreliability when using LLMs for annotation tasks without rigorous validation (Baumann et al. 2025). Given these concerns, we opt for full manual evaluation for all metrics except source category in this study.
@@
For the first step, we selected the queries related to baby care and pregnancy from the 10.4 million distinct Bing search queries in the Open Resource for Click Analysis in Search (ORCAS) (which is a public click-based dataset associated with information retrieval tasks [^1] This search query dataset contains no personally identifiable information and has been filtered using a k-anonymity threshold to exclude infrequent queries. Offensive and pornographic content has been removed, meaning some extreme queries may be absent.(Craswell et al. 2020). Using this dataset allows us to include real search queries actually utilized by the users when seeking information or advice on baby care and pregnancy-related questions. For the first filtering step, we used the keywords “baby” and “babies” for the topic baby care and “pregnant” and “pregnancy” for the topic pregnancy. For the second step, we filtered for queries with a minimum length of 3 words, as 2-word queries (e.g., “baby teeth,” “baby eyes”) tend to be too vague. Additionally, we removed almost identical queries. To do this, we applied tokenization, stemming, and stopwords-removal to the queries, and then removed the queries that were duplicates after this pre-processing step. As a result, we obtained a total of 15,500 queries.
In the third step, we conducted further filtering of the queries to ensure we include only those related to human baby care and/or pregnancy. This was necessary to filter out queries related to (1) baby care/pregnancy of an animal or plant [^2] We filtered out any plant-related queries that might use pregnancy-related terminology.; (2) information about pregnancy of a celebrity, fictional character, or a specific person; (3) a name, song, or item that simply contains the words “baby” or “pregnancy” but is not related to seeking advice or information on pregnancy. For this step, we relied on annotation using gpt-4o-mini. Following best practices for LLM-based annotation (Baumann et al. 2025), we validated the model’s performance on a subset of 50 manually annotated queries before applying it to our full dataset. The gpt-4o-mini classifications overlapped with our manual annotations 100% of the time. The exact prompt we used can be found in the Appendix. In total, after this step, we retained 9516 queries relevant to questions about human baby care and/or pregnancy.
@@
  • Numeric Mismatches: Different, but specific ranges, times, measurements or likelihoods that don’t overlap and can create confusion; any numerical or temporal or likelihood differences or mismatches, following a definition in prior work (Li, Raheja, and Kumar 2023);
@@ Results
Query sentiment is also significantly (Chi-Square test $p<0.001$) associated with the frequency of appearance of FS. FS appears nearly twice as often for negative (48.41%) compared to neutral (24.84%) queries. The difference is also statistically significant ($p=0.0265$) for AIO, with AIO appearing most often for negative (86.62%) and neutral (85.99%) queries, and slightly less often for positive ones (76.43%) (as shown in Figure 2) (b).
@@ Discussion
Acknowledgments
We thank Ronald E. Robertson, Kaicheng Yang, Piotr Sapieżyński, and Jeffrey Gleason for their valuable feedback on this work. We also thank everyone in the Social Computing Group at the University of Zurich and the anonymous reviewers for their helpful comments and suggestions.
The work of Aleksandra Urman, Elsa Lichtenegger, and Aniko Hannak was supported by Swiss National Science Foundation Grant 215354; and the work of Robin Forsberg was supported by the Kone Foundation.